OpenStudy (anonymous):

If the principal P = $700, the rate r = 6 1/2 %, and time t = 1 year, find the following. What is the amount of interest? What is the future value?

OpenStudy (anonymous):

Is this simple or compound interest?

OpenStudy (anonymous):

??

OpenStudy (dumbcow):

since its only for 1 year, and nothing is said about compounding, essentially its simple interest FV = 700(1+r) FV = 700 + 700(0.065) FV = 700 + 45.5 <--interest FV = 745.5

OpenStudy (anonymous):

thank you that helped me alot